OpenStudy (anonymous):

Solve the equation 1 over 3 x + 2 = 10 for x. 36 30 24 12

OpenStudy (ranga):

Does the equation look like this\[\Large \frac{ 1 }{ 3}x + 2 = 10\]?

OpenStudy (anonymous):

3x+2=10 x=8/3

OpenStudy (ranga):

x = 24

OpenStudy (anonymous):

1/3x=8 x=24